20g of impure calcium carbonate reacts with HCl and forms 6.6g of carbon dioxide. Calculate the purity of calcium carbonate.

Grade:10

2 Answers

KOUstubh
5 years ago
75 %
sd
write the equation
then balance it
calulate the mass of CO2that should be actually produced 
and then divid e the given amount of co2 by actual amount multiplied by 100
Abhiraj Pathak
5 years ago
CaCO3 + 2HCl → CO2 + H2O + CaCl2
1 mole of CaCO3 gives 1 mole CO2
i.e 100g of CaCO3 produces 44g of CO2
Thus 6.6g is produced by (100 x 6.6)/44 = 15g of CaCO3.
Therefore, % purity of CaCO3 = 15/20 = 0.75 = 75%